# SSL4RL: Revisiting Self-supervised Learning as Intrinsic Reward for Visual-Language Reasoning

### v1

@@ 4.1 Experiments on Vision-Language Reasoning Tasks
η Takeaway SSL as Intrinsic Reward Sharpens VLM Reasoning. The SSL4RL paradigm demonstrably enhances vision-language reasoning by repurposing SSL tasks as intrinsic rewards. It deepens the perception and understanding of the image itself, leading towards more precise visual attention and less language bias. Task Choice is Critical. SSL tasks show effectiveness when their inherent semantic aligns with core reasoning skills (e.g, Position and Rotation), while an inappropriate task may induce negative transfer and hinder downstream performance.
4.2 Ablation Study
4.2.1 Task Difficulty Scaling
The impact of task difficulty varies considerably across different SSL tasks, as evidenced by the results in Figure 4 (full results in Appendix C). The performances for the Contrastive task shows a marked improvement upon increasing its difficulty, elevated from 69.27% to 77.89% on MMBench and from 61.90% to 65.00% on SEED-Bench. Conversely, the benefits are marginal for the Position task and even detrimental for Rotation and Jigsaw. A plausible explanation is that the Contrastive task’s inherent simplicity as a binary discrimination problem means that raising the difficulty compels the model to learn more robust and informative features. On the other hand, exacerbating the difficulty of already challenging tasks like Jigsaw might induce over-specialization to the SSL objective, resulting in a negative transfer where the learned representations are less transferable or even counterproductive for downstream vision-language reasoning.
4.2.2 Model Size Scaling
To investigate the scalability of the SSL4RL paradigm, we apply it to the larger Qwen2.5-VL-7B-Instruct model. The results (Table 3 and Appendix Table 8) confirm that SSL4RL still yields improvements over the base model, notably enhancing Logical Reasoning on MMBench by 2.57% and Text Understanding on SEED-Bench by 3.57%. However, the gains are less pronounced than those observed with the 3B model. Even after increasing the task difficulty to a 5×5 grid for Jigsaw and Position tasks, no significant improvement was observed (see Appendix C).
We attribute these diminishing returns to a fundamental ceiling effect imposed by the predefined SSL tasks. The absolute difficulty of the four SSL tasks is fixed, presenting a suitable challenge for the 3B model but potentially failing to engage the full capabilities of a 7B model. The larger model’s enhanced abilities could render the tasks trivial, thereby weakening their effectiveness as a learning signal. This underscores a key insight: the efficacy of an SSL task is contingent on its ability to present a non-trivial challenge commensurate with the model’s capacity. Consequently, a primary challenge for future work is the design of adaptive or inherently more complex SSL objectives that can continue to provide a learning signal for large-scale models.
Base Qwen2.5-VL-7B 76.49 84.68 85.69 84.66 84.49 89.15 86.37
SSL4RL-7B Rotation 78.70 86.16 87.13 85.91 88.47 88.92 87.50
Jigsaw 74.87 83.26 85.18 85.24 86.28 89.09 86.27
Contrastive 77.73 83.19 84.58 85.89 85.29 87.74 86.25
Position 79.06 82.13 85.29 85.02 85.42 88.95 86.25
Maximal Improvement $\uparrow$ 2.57 $\uparrow$ 1.48 $\uparrow$ 1.44 $\uparrow$ 1.25 $\uparrow$ 3.98 $\downarrow$ 0.06 $\uparrow$ 1.13
4.2.3 Task Combination
Figure 5: Test accuracy of SSL tasks.
η Takeaway Goldilocks Principle of Task Difficulty. The effectiveness of an SSL task is contingent on its difficulty being appropriately matched to the model’s capacity. Insufficient challenge provides a weak learning signal, while excessive difficulty leads to negative transfer. Diminishing Returns with Model Size. The performance gains from the four SSL tasks diminish as model size increases (from 3B to 7B), suggesting designing SSL tasks with inherently higher complexity for large-scale models. Non-additivity of Rewards. A naive combination of multiple SSL rewards does not yield cumulative improvements, indicating potential optimization conflicts and underscoring the need for sophisticated integration strategies rather than simple averaging.
4.3 Experiments on Vision-Centric Tasks
@@
Benchmarks. We further evaluate the SSL4RL paradigm on vision-centric tasks using the ImageNet-1K dataset (Deng et al., 2009), which comprises approximately 1.3 million images across 1,000 categories. From this dataset, we construct a balanced subset of 100,000 training and 10,000 test images. To probe reasoning capabilities at varying difficulty levels, we design three question types: (1) Completion: directly answer with the species name; (2) Choice-20: select the correct species from 20 candidates; (3) Choice-200: select the correct species from 200 candidates.
@@ 5 Extension to Other Domains: An Example on Graph
Tasks and Reward Definitions. We introduce three graph-based SSL tasks, defined as follows: (1) Attribute Mask (Jin et al., 2020; Hu et al., 2019): A subset of node descriptions is randomly masked. The reward is quantified by the model’s accuracy in reconstructing the original masked features. (2) Neighbor Prediction (Kipf & Welling, 2016): For a target node within a partially observed graph, the model is rewarded for correctly identifying its adjacent nodes. (3) Link Prediction (Hu et al., 2020; Hou et al., 2022): Given a pair of nodes and a partial graph structure, the model receives a reward for accurately classifying the presence or absence of an edge connecting them.
Benchmarks. We evaluate our method on benchmark datasets curated from TAGLAS (Feng et al., 2024), a comprehensive collection of text-attributed graphs. The evaluation encompasses two key tasks: (1) Node-level classification on the Cora and PubMed co-citation graphs and the WikiCS page relation graph; (2) Link-level prediction on the Products co-purchase graph and the fb15k237 and wn18rr knowledge graphs.
Results and Observations. The results in Table 5 demonstrate the successful application of SSL4RL to graph-structured data. The 3B model shows substantial improvements, with gains up to 13.79% on average, while the 7B model exhibits diminishing returns, mirroring our observations in the visual domain and reinforcing the “difficulty-capacity matching” principle. Furthermore, the relative effectiveness of the self-supervised tasks is contingent upon the nature of the downstream objective. Tasks emphasizing structural reasoning (Link and Neighbor Prediction) yield better performance on relation-centric tasks such as link prediction. Conversely, tasks focused on feature reconstruction (Attribute Mask) demonstrate a comparative advantage on node classification benchmarks. These findings not only validate the generalizability of the SSL4RL framework beyond the visual modality but also highlight the critical importance of aligning the pretext task’s inductive bias with the target application.
Table 5: Test performance (%) on downstream graph tasks.
@@
Base model Qwen2.5-7B-Instruct 64.80 69.86 49.15 50.50 30.10 32.50 57.73
SSL4RL-7B Attribute 63.80 59.55 54.23 51.30 37.00 32.70 49.76
Neighbor 63.10 70.87 52.54 43.47 33.00 34.00 49.49
Link 67.70 72.27 55.93 48.89 21.30 39.50 50.93
@@ 6 Conclusions
We have introduced SSL4RL, a framework that repurposes self-supervised tasks as verifiable reinforcement learning rewards for post-training vision–language models. Our study shows that SSL4RL not only improves performance on vision-centric benchmarks such as ImageNet-1K, but also enhances multimodal reasoning, achieving substantial gains on MMBench and SEED-Bench. These findings suggest a broader principle: verifiable and scalable supervision signals are already embedded in self-supervision, and with proper task selection they can drive alignment and reasoning in VLMs without reliance on external verifiers, judges, or costly human labels. Looking forward, SSL4RL opens a path toward safer and more capable multimodal foundation models by unifying the strengths of self-supervision and reinforcement learning.
@@ Appendix A Detailed Results of MMBench
In this section, we present the detailed leaf task results of MMBench in Tables 6 and 7.
@@ Appendix B Results of SSL4RL 7B-model on SEED-Bench
Maximal Improvement $\uparrow$ 3.57 $\uparrow$ 2.71 $\uparrow$ 0.83 $\uparrow$ 0.88 $\uparrow$ 1.03 $\uparrow$ 1.78 $\uparrow$ 3.57 $\uparrow$ 2.13 $\uparrow$ 0.78 $\uparrow$ 0.86
@@ Appendix C Detailed Results of the Ablation Study about Difficulty
From Table 9 to Table 12, we provide detailed experimental results for the ablation study on task difficulty of MMBench and SEED-Bench.

### v2

@@ 4.1 Vision-Language Reasoning Tasks
Takeaway SSL as Intrinsic Reward Sharpens VLM Reasoning. The SSL4RL paradigm demonstrably enhances vision-language reasoning by repurposing SSL tasks as intrinsic rewards. It deepens the perception and understanding of the image itself, leading towards more precise visual attention and less language bias. Task Choice is Critical. SSL tasks show effectiveness when their inherent semantic aligns with core reasoning skills (e.g, Position and Rotation), while an inappropriate task may induce negative transfer and hinder downstream performance.
4.2 Open-ended Vision-Language Task: Image Captioning
SSL4RL requires no human labels, external verifiers, or heuristic judges, yet produces dense and scalable reinforcement signals. Intuitively, it has potential for open-ended tasks where ground truth is ill-defined and human annotations are expensive. To evaluate SSL4RL’s potential on open-ended tasks, we leverage a recent image captioning platform, CapArena [cheng2025caparena], where captions from a test model are compared against those from strong baseline models using GPT-4o as a judge, with human references provided for context. See Appendix I.2 for more details. As shown in Table 2. SSL4RL consistently improves over the base model, with the largest performance gain of 8.14 points (56.45 vs. 48.31). This clearly shows that the self-supervised rewards provide a meaningful learning signal even in the absence of a verifiable ground truth.
Qualitative Insights. We conduct a qualitative comparison with examples in Appendix Table 26, shed light on how SSL4RL enhances open-ended generation: (1) Enhanced Detail Capture. SSL4RL models correctly capture more details in the image. For instance, our SSL4RL model correctly identifies a ”scoreboard” displaying ”1 0 0” and a ”disabled persons’ sign,” which the base model either misinterprets or omits. (2) Improved Spatial Reasoning. SSL4RL models tend to use more precise spatial descriptors like ”behind the fence,” ”to the left of,” and ”the center of the court”, aligning well with the spatial reasoning ability required by SSL pretext tasks. In summary, by providing a dense, automated learning signal derived from the data’s intrinsic structure, our method successfully improves model performance on the complex, open-ended task of image captioning.
Table 2: Scores on CapArena Platform. GPT-Score: the score against GPT-4o. Cog-Score: the score against CogVLM-19B. CPM-Score: the score against MiniCPM-8B.
Category Model GPT-Score Cog-Score CPM-Score Average
Base Qwen2.5-VL-3B 0.00 6.48 92.96 48.31
SSL4RL-3B Rotation 19.15 49.30 96.48 55.28
Jigsaw 8.87 54.93 98.59 55.64
Contrastive 4.96 57.95 100.00 56.45
Position 0.00 59.15 92.25 51.04
4.3 Vision-Centric Tasks: ImageNet Classification
@@
Benchmarks. We further evaluate the SSL4RL paradigm on vision-centric tasks using the ImageNet-1K dataset [imagenet], which comprises approximately 1.3 million images across 1,000 categories. From this dataset, we construct a balanced subset of 100,000 training and 10,000 test images. To probe reasoning capabilities at varying difficulty levels, we design three question types: (1) Completion: directly answer with the species name; (2) Choice-20: select the correct species from 20 candidates; (3) Choice-200: select the correct species from 200 candidates.
@@
4.4 Robustness Analysis
Settings. To evaluate our model’s robustness, we conduct a comprehensive evaluation under various image perturbations. We designed two levels of perturbation—weak and strong—to assess the models’ resilience to visual corruptions. The weak perturbation applies a series of common image transformations, each with a probability of 0.5, including color jittering, random conversion to grayscale, Gaussian blur, and horizontal flipping. The strong perturbation builds upon the weak one by incorporating an additional multi-crop strategy. This strategy generates two 224×224 pixel crops per image via randomly resized cropping (scale range [0.08, 1.0]), presenting a more significant challenge to the model’s perception. Examples of these perturbed images are provided in Figure 6.
Results. As shown in Table 18 and Table 19, our method consistently demonstrates superior robustness across both perturbation levels. For instance, under weak perturbation, our SSL4RL-Rotation model achieves an average score of 72.11%, outperforming the base model at 66.80%. This performance gap is maintained under strong perturbation, where our SSL4RL-Position model scores 64.37% compared to the base model’s 59.62%. These results confirm that our SSL4RL strategy effectively enhances the model’s robustness to photometric variations.
4.5 Ablation Study
4.5.1 Data Volume Scaling
A key question for data-driven methods is the ability to leverage increasing amounts of data. To characterize the relationship between performance and data scaling for SSL4RL, we progressively expand our training dataset, obtaining three regimes of increasing data volumes: (1) Base Set: $\sim$4,000 samples from MMBench. (2) Extended Set: $\sim$18,000 samples from a mixture of MMBench and SEED-Bench. (3) Full Set: $\sim$118,000 samples from MMBench, SEED-Bench, and ImageNet.
We select the representative Position for evaluation. The results, detailed in Table 4, demonstrate a clear positive scaling relationship across almost all subtasks. When scaling from Base Set to Extended Set, the average performance on MMBench improved from 80.08% to 81.38%, a gain of +1.30%. A further expansion to the Full Set yielded an additional +1.04% improvement, reaching 82.42%. This consistent, monotonic improvement suggests that the effectiveness of SSL4RL continues to benefit from larger, more diverse datasets. We provide more discussions in Appendix E.
Table 4: Performance (%) on MMBench with increasing training data volumes.
Model Training Volumes Logical Relation Attribute Coarse Cross-Inst. Single-Inst. Average
Qwen2.5-VL-3B - 61.77 41.54 76.62 73.55 64.32 82.06 72.99
SSL4RL-Position 4,000 67.65 77.19 82.22 82.15 66.51 85.39 80.08
18,000 68.71 79.26 85.03 81.65 70.16 86.47 81.38
118,000 68.77 79.29 84.26 83.66 72.87 88.31 82.42
4.5.2 Model Size Scaling
To investigate the scalability of the SSL4RL paradigm, we apply it to the larger Qwen2.5-VL-7B-Instruct model. The results (Table 5 and Appendix Table 9) confirm that SSL4RL still yields improvements over the base model, notably enhancing Logical Reasoning on MMBench by 3.70% and Text Understanding on SEED-Bench by 3.57%. However, the gains are less pronounced than those observed with the 3B model. Even after increasing the task difficulty to a 5×5 grid for Jigsaw and Position tasks, no significant improvement was observed (see Appendix C).
We attribute these diminishing returns to a fundamental ceiling effect imposed by the predefined SSL tasks. The absolute difficulty of the four SSL tasks is fixed, presenting a suitable challenge for the 3B model but potentially failing to engage the full capabilities of a 7B model. The larger model’s enhanced abilities could render the tasks trivial, thereby weakening their effectiveness as a learning signal. This underscores a key insight: the efficacy of an SSL task is contingent on its ability to present a non-trivial challenge commensurate with the model’s capacity. Consequently, a primary challenge for future work is the design of adaptive or inherently more complex SSL objectives that can continue to provide a learning signal for large-scale models. In Appendix J, we provide a preliminary exploration of harder SSL4RL tasks on 7B models.
Base Qwen2.5-VL-7B 76.49 84.68 85.69 84.66 84.49 89.15 86.37
SSL4RL-7B Rotation 78.70 86.16 87.13 85.91 88.47 88.92 87.50
Jigsaw 80.19 84.64 88.02 85.70 84.53 90.93 87.73
Contrastive 77.73 83.19 84.58 85.89 85.29 87.74 86.25
Position 79.06 82.13 85.29 85.02 85.42 88.95 86.25
Maximal Improvement $\uparrow$ 3.70 $\uparrow$ 1.48 $\uparrow$ 2.33 $\uparrow$ 1.25 $\uparrow$ 3.98 $\uparrow$ 0.98 $\uparrow$ 1.36
4.5.3 Base Model Choice
To evaluate whether the benefits of SSL4RL extend to architectures beyond the Qwen2.5-VL series, we select Gemma3-4B [gemmateam2025gemma3technicalreport], a recent and powerful model from Google with a distinct architecture. To ensure a fair comparison under computational constraints, we adapted our training setup by halving the batch size to 256 while meticulously preserving all other hyperparameters and training procedures. Results presented in Appendix Table 22 provide strong evidence for the general applicability of our method. On the Gemma3 base model, our SSL4RL models yield a consistent and notable average performance improvement of 2.88% on MMBench. The gain is particularly pronounced in Cross-instance Perception (5.76%), Logical Reasoning (4.35%), and Attribute Reasoning (4.13%). Moreover, the relative efficacy of the individual SSL4RL tasks is remarkably consistent with our prior findings. As with Qwen2.5-VL, tasks like Rotation and Position confer the most significant benefits, while the simpler Contrastive task shows more modest gains. The consistent results validate that SSL4RL is a general-purpose principle for VLMs, not an artifact of a particular model family.
4.5.4 Task Difficulty Scaling
The impact of task difficulty varies considerably across different SSL tasks, as evidenced by the results in Figure 4 (full results in Appendix C). The performances for the Contrastive task show a marked improvement upon increasing its difficulty, elevated from 69.27% to 77.89% on MMBench and from 61.90% to 65.00% on SEED-Bench. Conversely, the benefits are marginal for the Position task and even detrimental for Rotation and Jigsaw. A plausible explanation is that the Contrastive task’s inherent simplicity as a binary discrimination problem means that raising the difficulty compels the model to learn more robust and informative features. On the other hand, exacerbating the difficulty of already challenging tasks like Jigsaw might induce over-specialization to the SSL objective, resulting in a negative transfer where the learned representations are less transferable or even counterproductive for downstream reasoning.
4.5.5 Task Combination
Figure 5: Test accuracy of SSL tasks.
Takeaway Goldilocks Principle of Task Difficulty. The effectiveness of an SSL task is contingent on its difficulty being appropriately matched to the model’s capacity. Insufficient challenge provides a weak learning signal, while excessive difficulty leads to negative transfer. Diminishing Returns with Model Size. The performance gains from the four SSL tasks diminish as model size increases (from 3B to 7B), suggesting designing SSL tasks with inherently higher complexity for large-scale models. Non-additivity of Rewards. A naive combination of multiple SSL rewards does not yield cumulative improvements, indicating potential optimization conflicts and underscoring the need for sophisticated integration strategies rather than simple averaging.
@@ 5 Extension to Other Domains: An Example on Graph
Having established SSL4RL for vision-language reasoning, we now explore its broader applicability. The paradigm’s core principle—generating rewards from data transformations—naturally extends to any domain with rich structural information. Beyond images, graph-structured data presents a compelling candidate, given its explicit relational semantics that are amenable to various pretext tasks. In this section, we empirically validate this potential by adapting SSL4RL to the graph domain. We provide details of tasks, reward definitions, and benchmarks in Appendix D. The results in Table 6 and Appendix Table 16 demonstrate the successful application of SSL4RL to graph-structured data. The 3B model shows substantial improvements, with gains up to 13.79% on average, while the 7B model exhibits diminishing returns, mirroring our observations in the visual domain and reinforcing the “difficulty-capacity matching” principle. Furthermore, the relative effectiveness of the self-supervised tasks is contingent upon the nature of the downstream objective. Tasks emphasizing structural reasoning (Link and Neighbor Prediction) yield better performance on relation-centric tasks such as link prediction. Conversely, tasks focused on feature reconstruction (Attribute Mask) demonstrate a comparative advantage on node classification benchmarks. These findings not only validate the generalizability of the SSL4RL framework beyond the visual modality but also highlight the critical importance of aligning the pretext task’s inductive bias with the target application.
Table 6: Test performance (%) of 3B models on downstream graph tasks.
@@ 6 Conclusions
We have introduced SSL4RL, a framework that repurposes self-supervised tasks as verifiable reinforcement learning rewards for post-training vision–language models. Our study shows that SSL4RL not only improves performance on vision-centric benchmarks such as ImageNet-1K, but also enhances multimodal reasoning, achieving substantial gains on MMBench and SEED-Bench. These findings suggest a broader principle: verifiable and scalable supervision signals are already embedded in self-supervision, and with proper task selection, they can drive alignment and reasoning in VLMs without reliance on external verifiers, judges, or costly human labels. Looking forward, SSL4RL opens a path toward safer and more capable multimodal foundation models by unifying the strengths of self-supervision and reinforcement learning.
@@
The Use of Large Language Models (LLMs)
In this work, LLMs are primarily employed for polishing the language of the manuscript to ensure grammatical correctness and coherence. Importantly, all conceptual development, theoretical analysis, experimental design, and result interpretation are conducted independently by the authors. The use of LLMs is strictly limited to auxiliary tasks, ensuring that the scientific contributions of this paper remain entirely unaffected by such tools.
@@ Appendix A Detailed Results of MMBench
In this section, we present the detailed leaf task results of MMBench in Tables 7 and 8.
@@ Appendix B Results of SSL4RL 7B-model on SEED-Bench
Maximal Improvement $\uparrow$ 3.57 $\uparrow$ 2.71 $\uparrow$ 0.83 $\uparrow$ 0.88 $\uparrow$ 1.03 $\uparrow$ 1.78 $\uparrow$ 1.13 $\uparrow$ 2.13 $\uparrow$ 0.78 $\uparrow$ 0.86
@@ Appendix C Detailed Results of the Ablation Study about Difficulty
From Table 10 to Table 13, we provide detailed experimental results for the ablation study on task difficulty of MMBench, SEED-Bench, and ImageNet1k.
@@
5x5 97.73 72.15 93.75 75.24 90.07 95.83 70.00 67.80 98.53 63.33 95.81 90.07 90.71 91.62 67.02 91.86 98.23 95.28 95.51 81.00 86.08
Table 13: Test performance (%) of models trained with different task difficulties on ImageNet-1K.
Model Difficulty Completion Choice10 Choice200
Qwen2.5-VL-3B - 24.93 85.22 57.10
Position 2x2 28.76 92.35 67.14
3x3 27.37 88.99 59.93
Contrastive Weak 26.84 89.51 61.78
Strong 26.93 89.44 61.31
Rotation 90-degree 29.19 87.26 58.48
45-degree 29.91 89.94 60.52